if Cu2+/Cu electrode is decreased 10 times then what will be the chamge in reduction potentiol?

Grade:12

1 Answers

Shailendra Kumar Sharma
6 years ago
Cu+2 +2e → Cu

Eold=Eo + (0.0591/2) log (Cu+2)
with the decrease in concentration by Cu+2/10
Enew=Eo + (0.0591/2) log (Cu+2/10)
       = Eo + (0.0591/2) log (Cu+2)-(0.0591/2)log10
       =Eold – 0.0295